A student wanted to synthesize 4-(5-methylhex-3-yn-1-yl)phenol. The student\'s key step was to be the SN2 reaction of an acetylide with 4-(2-bromoethyl)phenol. However, the student isolated none of the desired product. What went wrong? Draw the two products the student isolated instead. Assume a 1:1 mixture of reagents.

A student wanted to synthesize 4-(5-methylhex-3-yn
Answer #1
Concepts and reason

A reaction product is given as, 4(5methylhex3yn1yl)phenol{\rm{4 - }}\left( {{\rm{5 - methylhex - 3 - yn - 1 - yl}}} \right){\rm{phenol}} whose reactants are proposed as 4(2bromoethyl)phenol{\rm{4 - }}\left( {{\rm{2 - bromoethyl}}} \right){\rm{phenol}} and acetylide but expected product is not obtained in the reaction and other products are formed which needs to be formed.

The mechanism expected to be SN2{{\rm{S}}_{\rm{N}}}{\rm{2}} , therefore, the reaction will occur in the presence of strong nucleophile with haloalkane and nucleophile will be substituted at the carbon containing halogen.

Fundamentals

SN2{{\rm{S}}_{\rm{N}}}2 reaction is bimolecular nucleophilic substitution reaction which takes place between a haloalkane and a nucleophile. The reaction takes place in one step through a transition state and the reaction is stereospecific in nature. Consider the following scheme of a SN2{{\rm{S}}_{\rm{N}}}2 reaction:

Н
Н
Nu-
Nu
Nu
R
Nu Nucleophile|
|
RI, R2 H, alkyl group

In the reaction, the nucleophile attacks from the back side of the reaction as the front side is hindered by the leaving group.

Consider the given reactant and the reagent used to form the expected product as shown below:

-c=c
+
HO
3-methylbut-1-yn-1-ide
4-(2-bromoethyl)phenol
(expexted product) (5-methylhex-3-yn-1-yl)benzene

But acetylide also act as a strong base and can abstract the acidic proton as well. Therefore, the choice of nucleophile went wrong.

In the given reaction nucleophile used is an acetylide which will abstract the phenolic proton and undergo acid base reaction to form the following product:

Br
-cc
НО
3-methylbut-1-yn-1-ide
4-(2-bromoethy)phenol
-Br
-CECH
3-methylbut-1-yne
4-(2-bromoethyl)phenolate
(Products)

Ans:

Hence, the product obtained by the given reactants is as followed:

.Br
-cEc:
+
НО
3-methylbut-1-yn-1-ide
4-(2-bromoethyl)phenol
.Br
-CCH
+
3-methylbut-1-yne
4-(2-bromoethyl)phenolate
(Products